OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)?

OpenStudy (texaschic101):

First we find the slope using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (4,0) x1 = 4 and y1 = 0 (2,6) x2 = 2 and y2 = 6 now we sub slope(m) = (6 - 0) / (2 - 4) slope(m) = -6/2 which reduces to -3 so now we know that the slope is -3. Now we can use our point slope form : y - y1 = m(x - x1) slope(m) = -3 use either of your set of points...(4,0) x1 = 4 and y1 = 0 y - 0 = -3(x - 4) === >point slope form or if you used the other points (2,-4) y - (-4) = -3(x - 2) y + 4 = -3(x + 2) ===> point slope form